# Large Language Models Play StarCraft II: Benchmarks and A Chain of Summarization Approach

### v1

@@ 1 Introduction
    We developed the TextStarCraft II environment, providing a platform for natural language interaction with StarCraft II.
@@
    We designed the Chain of Summarization (CoS) method, enhancing the language model’s multi-layer reasoning capabilities and global decision-making speed, defeating the built-in AI with a harder difficulty level in the latest game versions and maps.
@@
    The emergence of advanced tactics, including reconnaissance (Fig 5.a and Fig 5.b), economic expansion, defensive building construction (Fig 5.c), research upgrades (Fig5.d), and unit transitions (Fig 6.a, Fig 6.c, and Fig 6.d).Our experiments show that by utilizing LLMs without additional training, our AI achieves competitive performance in StarCraft II, effectively integrating human strategies. Moreover, our method adapts to the latest game versions and maps through adjustments in both prompts and scripts, showcasing its versatility and prompt-script based adaptability.
@@
    We invited 30 grandmaster-level StarCraft II players for blind assessments, comparing the knowledge mastery and strategy levels of various large models such as GPT3.5, GPT4, Claude2 and Bard.
2 Related work
StarCraft II Full Game AI: In the era of StarCraft I, a large number of researchers began to study StarCraft AI. BiCNet[16] approach significantly enhanced multi-agent coordination in StarCraft I, marking a key advancement in the field.With the introduction of the PySC2666https://github.com/google-deepmind/pysc2[23] learning environment for StarCraft II by DeepMind and the provision of tens of thousands of replays by Blizzard Entertainment, research on StarCraft II AI began to thrive. By 2019, DeepMind introduced AlphaStar[22], which employed a federated learning approach. It first defeated European professional players TLO and MANA. Later, it was tested on the European ladder server, reaching the level of mid-tier Grandmaster among human players. Ultimately, at the 2019 BlizzCon, it defeated the world’s top Zerg player, Serral [^7] https://www.youtube.com/watch?v=nbiVbd_CEIA. This achievement demonstrated that reinforcement learning methods can reach human top-tier levels in such a complex and changeable POMDP environment. This sparked researchers’ enthusiasm for StarCraft II, with subsequent works taking AlphaStar as a starting point. PySC2 has since become the only environment for StarCraft II Full Game research.After AlphaStar, research in StarCraft II global game AI can be categorized into two types based on whether it can achieve real-time human-computer gameplay.
Research Without Real-Time Human-AI Interaction
Mini-AlphaStar[12] aims to reduce the number of input variables without significantly compromising learning effectiveness. TG[10] achieves efficient StarCraft reinforcement learning using abstract forward models and transfer learning. HierNet-SC2[11] employed data mining combined with hierarchical reinforcement learning, skipping the supervised learning pre-training step and starting training the agent from scratch. DeepMind launched their latest creation, AlphaStar Unplugged[13], which used a large number of human player replays for offline reinforcement learning, and ultimately succeeded in defeating the previous AlphaStar agent trained through behavior cloning.
Research Involving Real-Time Human-AI Interaction
TStarBotsX[7] developed a new federated learning approach and utilized rule-guided policy search methods to aid in training the AI agent, successfully defeating master and grandmaster level players from China and the USA.SCC[26] almost perfectly replicates AlphaStar’s results and defeated the Chinese champion "Time" in human-computer combat. However, some players and researchers believe that during training and combat, the AI was not limited by reaction speed, EPM, etc. "DI-star" [^8] https://github.com/opendilab/DI-star In addition, DI-star also offers a version that can be locally deployed for home computers, allowing ordinary players to compete against it.In addition, DI-star also offers a version that can be locally deployed for home computers, allowing ordinary players to compete against it.
LLM Agent: Since the emergence of GPT3.5[14], research on LLM agents has begun to rise. Starting with the earliest works like React[31] and Autogpt [^9] https://github.com/Significant-Gravitas/AutoGPT, numerous studies have been inspired by this approach. Research utilizing the innovative environment, MineDojo [^10] https://github.com/MineDojo/MineDojo[6], has opened new avenues in AI research, focusing on extensive task diversity and open-ended world development, distinguishing it from previous reinforcement learning environments.GITM (Ghost in the Minecraft)[34]successfully decomposed tasks into subtasks and unlocked the entire Minecraft task tree by setting up a Decomposer, Planner, and Interface. Voyager[24] implemented the discovery of new items, unlocking skill trees, and exploring the unknown world through automatic curricula, skill libraries, and iterative prompting. In multi-agent scenarios, "Camel"[9] designed a role-playing framework, allowing language models to collaborate and align with human intent. "MetaGPT"[8] integrated SOPs (Standard Operating Procedures) into multiple agents, enabling LLM agents to collaborate in a pipeline of roles, accomplishing software development tasks. In social simulation, "Generative Agents"[15] innovatively built a virtual town using LLM agents, supporting 25 agents with distinct personalities to communicate and establish social relationships.
Although these studies about starcraft II achieved powerful agents, some could defeat StarCraft II’s Level 10 built-in AI (e.g., TG), while others could beat human professional players (e.g., "SCC", "DI-star"). However, all have shortcomings such as a lack of long-term strategies, unexplainable tactics, and the inability to adapt troop compositions based on the opponent’s status.From the current work on LLM agents, we see their potential in addressing issues like the lack of long-term strategies in StarCraft II Full Game. Through our research, we found that the mainstream LLM agent studies have chosen GPT3.5-turbo and GPT4 and achieved excellent results. Therefore, we chose the GPT3.5-turbo-16k model for our LLM. And the existing StarCraft II learning environment, PySC2, does not support natural language interaction. Hence, we proposed "TextStarCraft II", a StarCraft II environment that allows natural language interactions
3 TextStarCraft II
We introduce "TextStarCraft II", a novel text-based interface for the StarCraft II environment. This interface is specifically designed to enable LLM agents to interact in a way similar to that of Reinforcement Learning (RL) agents. Built upon the ’python-sc2’ framework, known for its extensive API and thorough documentation, TextStarCraft II provides detailed textual game data, which enhances the decision-making capabilities of LLM agents in the context of StarCraft II."
The environment is consist of original StarCraft II game client and two adapters (We developed):
  1. 1.
    Observation-to-Text Adapter:Transforms game details into textual data, making the game’s complex scenarios comprehensible to LLM agents.
  2. 2.
    Text-to-Action Adapter:Interprets textual commands from the agents, turning them into actual game moves.
POMDP Formulation for TextStarCraft II
We formulate the TextStarCraft II as a multi-agent Partially Observable Markov Decision Process (POMDP). Denote the POMDP as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{R},\Omega,\mathcal{T},\mathcal{O}\rangle$. $\mathcal{S}$ is the state space of the underlying StarCraft II game engine, which is not directly observable to the agents. $\Omega$ is the observation space, in which an observation $o\in\Omega$ consists of resources, buildings, enemy information, etc. The observation $o\in\mathcal{V}^{K}$ is represented by natural language sentences, where $\mathcal{V}$ is the vocabulary with finite words in it. $\mathcal{A}=\{\mathcal{A}_{P},\mathcal{A}_{M}\}$ is the action space, which combines the primitive action space $\mathcal{A}_{P}$ of StarCraft II and the high-level action space $\mathcal{A}_{M}$ of the macro actions. Each macro action is a sequence of primitive actions to complete a subgoal, such as producing units, building structures, and researching technologies. $\mathcal{R}$ is the reward function, and the reward $\mathcal{R}(s,a)$ belongs to $\{-1,0,1\}$ representing lose, draw and win respectively. $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the deterministic transition function, which is a mapping to the next state with the current state and action given. The function $\mathcal{O}:\mathcal{S}\rightarrow\Omega$ maps from the state space to the observation space.
4 Chain of Summarization
Figure 1: Interacting with Large Language Models: The Enhanced Chain of Summarization Method in TextStarCraft II. A streamlined approach to LLM-driven strategic gameplay, the method starts with Initialization, converting the game’s initial data to text then queuing it for processing. Single-Frame and Multi-Frame Summarization follows, refining observations and summarizing them into actionable insights through advanced LLM reasoning. The Directive Formulation and Action Scheduling phase segments the insights into specific actions, populating an action queue for subsequent execution. Finally, Action Retrieval and Execution implements these actions within the game, closing the loop by converting new data into text for continuous interaction. This cyclical process advances LLM performance in the complex, strategic environment of TextStarCraft II.
Algorithm 1 Chain of Summarization (CoS)
1: The game round is current at the $i$-th period. Frame number $K$ per period. The observations of the $i-1$ period: $\{o_{(i-1)*K+1},\ldots,o_{i*K}\}$.
2: $\{a_{i*K}+1,\ldots,a_{(i+1)*K}\}$: predicted action sequence for the $i$-th round
3:Initialize observation summary list $\hat{O}=\{\}$
4:for $j=(i-1)*K+1$ to $i*K$ do
5:     Perform single-frame summarization as Eq. (1): $\hat{o}_{j}\leftarrow\mathcal{F}(o_{j})$
6:     Add $\hat{o}_{j}$ to summary list: $\hat{O}=\hat{O}\cup\{\hat{o}_{j}\}$
7:end for
8:Perform multi-frame summarization as Eq. (2): $\mathcal{S}_{i}=\mathcal{A}(\hat{o}_{(i-1)*K+1}\cdots\hat{o}_{i*K},\text{prompts})$
9:Apply Chain of Thought reasoning as Eq. (3): $\upsilon_{i}=\text{CoT}(\mathcal{S}_{i},\text{prompts})$
10:Extract the predicted $K$ step actions Eq. (4): $\{a_{i*K+1},\ldots,a_{(i+1)*K}\}\leftarrow\mathcal{E}(\upsilon_{i})$
11:return $\{a_{i*K+1},\ldots,a_{(i+1)*K}\}$
In the intensely competitive and highly complex game of StarCraft II, making efficient and professional decisions is crucial. Existing Chains of Thought (CoT) methods and their derivatives perform well in simple tasks but still face challenges in complex decision-making and comprehensive analysis required in StarCraft II. These include difficulties in understanding the game’s situation, slow interaction speed, and less professional decision-making.To address these challenges, inspired by the cache mechanism of computer hardware and frame skipping techniques in reinforcement learning, we propose the Chain of Summarization method. We optimize the decision-making process through the following key strategies:
Information Compression: We summarize and abstract game data, eliminating unnecessary informational noise to extract the most critical game information.
Inference Acceleration: Abandoning traditional single-frame reasoning, we employ multi-frame information for a comprehensive analysis of the game situation and strategy formulation. This approach significantly reduces the time from analysis to action implementation.
Global Understanding: By compressing key information and using multi-frame analysis to discern game trends, language models can understand complex game strategies and dynamics, enabling professional decision-making and promoting higher-level strategic planning.
The Chain of Summarization method is capable of handling the intricacies and fast-paced strategies of StarCraft II, facilitating efficient and strategically significant decision-making.
4.1 Components of Chain of Summarization
The Chain of Summarization comprises: Single Frame Summarization, Multi-Frame Summarization, and Action Extraction for Action Queue.
Single-frame Summarization (Level-1 Summarization):
Because it is redundant and difficult for LLMs to understand the raw observation representation from TextStarCraft.This method compresses and extracts essential information from the TextStarCraft‘s raw observation. During this phase, dense and intricate text-based observational data from TextStarCraft II is condensed, preparing it for more efficient subsequent processing. Denote the Single-frame Summarization module as $\mathcal{F}(.)$, the processed observation as $\hat{o}$, and the whole process is formally written as:
$\hat{o}=\mathcal{F}(o).$ (1)
The compression can be implemented in two different ways: a language model-based approach and a rule-based approach. The language model-based approach employs few-shot learning, which, although time-consuming, allows for an analysis more in line with game rules using Large Language Models (LLMs), resulting in better compression outcomes. To reduce processing time, we improved upon previous methods and developed a rule-based approach. This method employs rules for extraction and filtering to ultimately distill key information, providing a faster but less precise alternative. For the sake of accelerating interactions, our experiments primarily utilize the rule-based approach.
Multi-frame Summarization(Level-2 Summarization):
Existing works [32; 3] adapt the paradigm of querying LLMs at each time step for actions to solve decision-making tasks. However, due to the high computation cost for LLM inference and the long time duration of StarCraft game, the per-time query paradigm is not suitable. Additionally,LLMs need to understand game’s situation and evolving dynamics by analyzing multiple game frames simultaneously, facilitating complex reasoning.To address this, we introduce the Multi-frame summarization method, inspired by the caching mechanisms of computer hardware and the frame skipping techniques in reinforcement learning. This method aims to bridge the speed disparity between the fast-paced game environment and the slower LLM inference while ensuring real-time decision efficiency and enhance the LLMs understand ability in complex environment. Instead of querying LLMs at each time step, our method uses LLMs to aggregate the observations for $K$ steps and generate $K$ actions next corresponding to each time step by querying LLMs one time. Every $K$ frames (time steps), the level-1 summarized observations $\{\hat{o}_{(i-1)*K+1},\cdots,\hat{o}_{i*K}\}$ are aggregated into a period summarization,we call that situation overview, $\mathcal{S}_{i}$ for the period $i$. Formally, the aggregation process is written by:
$\mathcal{S}_{i}=\mathcal{A}(\hat{o}_{(i-1)*K+1}\cdots\hat{o}_{i*K},prompts)$ (2)
With period summarization, the CoT reasoning is then performed to process the complex reasoning for decision-making. In the TextStarCraft II game, the CoT process performs below reasoning steps by step with deliberately designed prompts:
  1. 1.
    Situation Overview: Review the current batch of L1 observations and summarize the prevailing state.
  2. 2.
    Situation Analysis: Evaluate the current state, gauge the strengths and weaknesses of both sides, and discern the strategies being deployed by each.
  3. 3.
    Strategic Planning: Based on the current situation, propose tactics that our side should adopt.
  4. 4.
    Opponent Strategy Analysis:Based on the current situation, we analyze the enemy’s strategy and intentions.
  5. 5.
    Suggestion Formulation: Given our strategic position and tactics, offer logical recommendations.
  6. 6.
    Decision-Making Phase: Formulate decisions based on insights and recommendations from the preceding five stages.
Assume $\upsilon_{i}$ as the thoughts for observation summarization $\mathcal{S}_{i}$ generated by the CoT process, written as:
$\upsilon_{i}=\text{CoT}(\mathcal{S}_{i},\text{prompts})$ (3)
Action Extraction for Action Queue
The action queue is a bridge connecting the multi-frame summarization results, denoted as $\upsilon_{i}$, with the TextStarCraft II environment, ensuring seamless communication between the LLM and TextStarCraft II. Within $\upsilon_{i}$, three components exist: analysis, suggestions, and decisions. To extract actionable items from the decisions, we utilize regular expression matching and similarity searching techniques through the action extractor $\mathcal{E}$. Extracted actions are then stored in the action queue, primed for execution within TextStarCraft II. Specifically, actions designated for the time steps between $i*K+1$ and $(i+1)*K$ are derived from thoughts in $\upsilon_{i}$. The extraction can be represented as:
${a_{i*K+1}\cdots a_{(i+1)*K}}=\mathcal{E}(\upsilon_{i}).$ (4)
4.2 Interacting with the Environment Using Chain of Summarization
Algorithm 2 Chain of Summarization Interaction in TextStarCraft II
1:Environment: env (TextStarCraft II); Chain length: K; Frame Queue: FQ; Action Queue: AQ; Chain of Summarization: CoS
2:Initialize the TextStarCraft II environment
[…]
4:$a_{0},...,a_{k}\leftarrow\text{CoS}(o_{0})$triangleright Generate initial actions
[…]
8:     $reward,obs\leftarrow\text{env.step}(action)$
[…]
10:     if $\text{len}(\text{FQ})\geq\text{K}$ then
[…]
12:         $a_{i*K},\ldots,a_{(i+1)*K}\leftarrow\text{CoS}(o_{(i-1)*K},\ldots,o_{i*K})$
[…]
14:     end if
15:end while
16:return total_reward
Figure 1 illustrates the interaction process within the TextStarCraft II environment using the Chain of Summarization method. Implementing the Chain of Summarization within game interactions involves four stages: initialization, single-frame and multi-frame summarization, directive formulation and action scheduling, and action retrieval and execution. The Chain of Summarization method excels in eliciting the macro-decision capabilities of the Large Language Model (LLM), enabling the generation of high-level, quality decisions from a plethora of game observations. Additionally, to further enhance interaction speed, we have set the environment’s interaction frequency to a 2:1 ratio, utilizing one frame of information in TextStarCraft II to represent two actual frames in the StarCraft II game, thereby interacting every two frames.
Stage 1: Initialization
Central to our method is the parameter k, which sets the number of text-based observations to be synthesized. TextStarCraft II converts the initial first-frame information into textual observations through an observation-to-text adapter, storing it in the frame queue,FQ. The Chain of Summarization then processes only one frame observation,and generates an action sequence of length k, ensuring the smooth progression of subsequent processes.
Stage 2: Single-Frame and Multi-Frame Summarization
During the core interaction stage, the system extracts K observations from the frame queue,FQ, specifically $o_{(i-1)*K}\cdots o_{i*K}$. These observations undergo an initial refinement through the Single-frame Summarization function $\mathcal{F}(.)$, resulting in a set of processed observations: $\hat{o}{(i-1)*K}\cdots\hat{o}{i*K}$. Following this, the Multi-frame Summarization phase commences, consisting of two key procedures. First, the aggregation function $\mathcal{A}$ consolidates the processed observations $\hat{o}_{(i-1)*K}\cdots\hat{o}_{i*K}$ into an integrated situation overview, designated as $\mathcal{S}_{i}$. Next, the system applies the Chain of Thought (CoT) reasoning method, CoT, directly to $\mathcal{S}_{i}$. This application synthesizes the data into a final, actionable output, represented as $\upsilon_{i}$, ready for execution. This streamlined process ensures efficient decision-making, drawing from comprehensive data analysis and strategic reasoning.
Stage 3: Directive Formulation and Action Scheduling
This stage refines the outputs from the Multi-frame Summarization $\upsilon_{i}$, segmenting them into distinct categories: analysis, suggestions, and decisions. Utilizing an action extractor $\mathcal{E}$, the system translates these comprehensive thought processes $\upsilon_{i}$ into a series of specific actions, denoted as ${a_{i*K}\cdots a_{(i+1)*K}}$. These formulated actions are then systematically stored in the action queue, referred to as AQ, awaiting deployment.
Stage 4: Action Retrieval and Execution
During each engagement cycle with TextStarCraft II, the system retrieves the next sequence of actions from the Action Queue AQ, translating them into executable game commands through a text-to-action adapter. After these commands are executed within the game environment, TextStarCraft II reciprocates by converting the new game data into textual format via an observation-to-text adapter, subsequently enqueuing it. Beyond the initial setup, every operational cycle of the Chain of Summarization method enriches the frame queue with an additional k frames of observational data $o$, perpetuating the data processing and decision-making continuum.
Our Chain of Summarization method, invoked once every k frames, demonstrates its efficiency in handling and responding to high-risk, multifaceted information in fast-paced, strategy-intensive environments. It underscores the effectiveness of the Chain of Summarization method in fostering rapid and prudent decisions in complex tasks, especially where traditional modes of thinking fall short. Through this innovative approach, we have significantly enhanced the language model’s performance in intricate environments like StarCraft II, setting the stage for potential future applications.
5 Experiment
5.1 Evaluation of the Chain of Summarization Method
Table 1: Performance Comparison Between Methods With and Without Chain of Summarization: Interaction frequency is approximately 55 times in 10 seconds, assuming each API call takes about 1.66 seconds and game . Using the Chain of Summarization significantly accelerates decision-making speed and reduces the number of API calls.
Method Number of API Calls Time Cost (hours)
With Chain of Summarization 700 7
Without Chain of Summarization 7,000 70
We chose GPT3.5-turbo-16k as our Large Language Model (LLM) and conducted tests in scenarios both with and without the Chain of Summarization (CoS) method. Notably, the python-sc2 package facilitates an interaction frequency of around 55 times every 10 seconds. Table 1 presents the comparative results of using CoS and not using it.
Without Multi-level Summarization: From an analysis of the number of API calls, playing a game of StarCraft II with game time 21 minutes,required approximately 7,000 API calls, leading to high costs and a total duration of roughly 70 hours(assuming that each API call requires approximately 1.66 seconds). In terms of decision-making, without the use of the chain of summarization method, the LLM lacked the capability for strategic planning and decision-making.
With Chain of Summarization: Analyzing the number of API calls, playing a game of StarCraft II necessitated around 700 API calls, resulting in relatively lower costs and an average game duration of about 7 hours. In terms of decision-making capabilities, the use of the chain of summarization allowed the LLM to comprehend complex information in StarCraft II, distill critical game data, formulate strategic overviews, and make effective decisions.
The Chain of Summarization not only significantly reduced the need for excessive API calls, saving time, but also notably enhanced the LLM’s game comprehension and strategic abilities. This enabled the LLM to analyze, judge, and strategize at a strategic level for complex scenarios like StarCraft II.

### v2

@@ A.1 Gameplay Overview
    Resource Management: Players gather two primary resources, minerals and vespene gas, to build structures and units.
@@
    Base Constructiont: Efficiently building and expanding bases is crucial for success. This includes creating resource collection structures, unit production facilities, and defensive fortifications.
@@
    Army Composition and Control: Players must construct a balanced mix of units and strategically control them during skirmishes and battles to defeat opponents.
@@
    Strategic Planning:The game requires foresight and the ability to adapt strategies in response to opponent moves.
A.2 E-Sports and Competitive Play
StarCraft II has been a cornerstone of the e-sports community since its launch in 2010. The game is renowned for its demanding mechanics and deep strategic complexity, which have made it a favorite among competitive players and spectators alike. Major tournaments such as the StarCraft II World Championship Series (WCS) and the Intel Extreme Masters (IEM) have featured the game, drawing competitors and viewers from around the globe.
Professional Scene: The professional SC2 scene is characterized by its high level of play and international participation. Players from South Korea, widely regarded as the epicenter of StarCraft competitive play, have traditionally dominated the global tournaments, although strong competitors from Europe, North America, and other parts of Asia have also made significant impacts.
Community and Legacy: StarCraft II’s legacy is supported by a vibrant community of players, fans, and content creators. The game’s strategic depth has also inspired numerous discussions, tutorials, and strategy guides, making it not just a game but a continual subject of study and analysis.
Appendix B Experiment Setup and Metrics Explanation
B.1 Experiment Setup
Agent and Opponent Selection: To ensure a consistent and controlled experimental environment, LLM agents are configured to play as Protoss against Zerg AI opponents. This setup allows for a systematic evaluation of strategic capabilities across a range of difficulties. The difficulty levels are defined as follows to cover a comprehensive spectrum of challenges:
Level 1 2 3 4 5
BLZ Difficulty Very Easy Easy Medium Hard Harder
Level 6 7 8 9 10
BLZ Difficulty Very Hard Elite Cheat Vision Cheat Money Cheat Insane
Table 5: AI Difficulty Levels for StarCraft II
Map Selection: For our experiments, we selected the Altitude LE and Ancient Cistern LE maps from the 2023 StarCraft II esports 1v1 ladder. These maps provide a variety of strategic challenges that are representative of current competitive play. Additional details about these maps can be found on their respective Liquipedia pages: https://liquipedia.net/starcraft2/Altitude_LE and https://liquipedia.net/starcraft2/Ancient_Cistern_LE.
Parameter Settings: The temperature parameter is fixed at 0.1 to focus on strategy-driven actions over randomness.
Game Version: Our experiments were conducted across three different versions of the game to ensure robustness and applicability of the results. The versions tested include Patch 5.0.11, Patch 5.0.12, and Patch 5.0.13.
B.2 Evaluation Metrics:
Building on StarCraft II’s established player performance analytics, our evaluation framework in TextStarCraft II integrates these insights with custom adaptations to comprehensively assess LLM agents’ gameplay strategies.
Win Rate: The most crucial metric, Win Rate directly reflects the agent’s capability and performance within the environment. It is calculated as the percentage of games won by the agent out of the total games played.
Population Block Ratio (PBR): PBR is a measure of the agent’s macro-management effectiveness, particularly in resource allocation and population growth. It is defined as:
PBR=Time at Population CapGame $\text{PBR}=\frac{\text{Time at Population Cap}}{\text{Game Duration}}$ (5)
Here, the metric calculates the ratio of time spent at population capacity to the total game time until the agent first reaches maximum supply (200/200). A higher PBR indicates poorer macro-strategic decision-making and ineffective planning.
Resource Utilization Ratio (RUR): RUR evaluates the agent’s efficiency in resource management throughout the game. It is calculated as:
RUR=Total Minerals + Total Gas UsedGame $\text{RUR}=\frac{\text{Total Minerals + Total Gas Used}}{\text{Game Duration}}$ (6)
This metric assesses the total resources used against the duration of the game until the agent first reaches maximum supply. A higher RUR signifies underutilization of resources, reflecting weaker macro-strategic capabilities.
Average Population Utilization (APU): APU measures the efficiency in utilizing the available population capacity. It is calculated as:
$\text{APU}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\text{Used Population at }i^{th}\text{ step}}{\text{Population Cap at }i^{th}\text{ step}}\right)$ (7)
The metric averages the ratio of used population to population cap until the agent first reaches maximum supply. A higher APU suggests better utilization of population and more effective macro-management.
Technology Rate (TR): TR assesses the extent to which the agent explores and utilizes the technology tree. It is defined as:
TR=Completed TechnologiesTotal Technologies $\text{TR}=\frac{\text{Completed Technologies}}{\text{Total Technologies Available}}$ (8)
This metric calculates the proportion of completed technologies and buildings to the total available, from the start to the end of the game. TR indicates the agent’s tendency towards technological advancement and is not necessarily indicative of better or worse performance.
Data Collection and Analysis: Due to technical constraints with the sc2reader library for AI-generated replays, we utilized gameplay log data for in-depth analysis. This approach ensures a thorough and precise evaluation based on our tailored metrics.
B.3 Compute Resource
Training Phase: To fine-tune open-source Large Language Models, we utilized two NVIDIA A100 40GB GPUs, providing the necessary computational power for extensive model training. The entire experiment took approximately 70 hours.
Testing Phase: The development and testing of these fine-tuned models were performed using two NVIDIA A100 40GB GPUs. For running the StarCraft II environment, we employed an NVIDIA GeForce RTX 3060 GPU paired with a 13th Gen Intel(R) Core(TM) i5-13400F processor operating at 2.50 GHz, ensuring smooth execution of the game simulations.
B.4 Human Experts
In the course of our research, we engaged 30 StarCraft II masters and grandmasters, including professional and semi-professional players, to participate in our experiments. To safeguard their privacy, we have chosen not to disclose their personal information in this paper.
Each participating player was compensated with 15 USD for their involvement in the study.
Appendix C Experiment Figure
Figure 5: Comparison of Unit Preferences: Analysis of LLM Agent in 30 Games vs. Professional Players in Esports Tournaments from 2015, 2020, and 2023
Table 6: Statistical Evaluation of Models
Mean Score Std Dev Pearson Corr T-test p-value Levene p-value
Model GPT-4 Human GPT-4 Human
Bard 3.70 3.79 1.02 1.12 0.00 0.4078 0.4718
ChatGPT 5.00 4.42 0.83 0.97 0.00 0.0000 0.0001
Claude2 4.69 3.91 0.70 1.12 -0.00 0.0000 0.0000
GPT4 5.17 4.77 0.71 0.93 -0.00 0.0000 0.0001